OpenStudy (anonymous):

can someone show me how to complete this problems.. 1 ) 4+2/x all over x/4+1/8 2) square root of 6a - 4 square root of 54a - 4 square root of 216a thanks in advance

OpenStudy (aum):

Is this the problem? \[\Large \frac{\frac{4+2}{x}}{\frac{x}{4 + \frac 18}} \]

OpenStudy (aum):

If you post a screenshot of the questions it will be easier to help you with the problems.

OpenStudy (anonymous):

yes that is the problem

OpenStudy (aum):

\[ \Large \frac{\frac{4+2}{x}}{\frac{x}{4 + \frac 18}} = \large \frac{6}{x} * \frac{4 + \frac 18}{x} = \frac{6}{x} * \frac{33}{8x} = \frac{3}{x} * \frac{33}{4x} = \frac{99}{4x^2} \]But I don't think that is the problem even though you verified it is. That is why a screenshot would help. I think the problem may be: \[ \Large \frac{4 + \frac 2x}{\frac x4 + \frac 18} = \frac{\frac{4x + 2}{x}}{{\frac{2x+1}8} } = \large \frac{4x+2}{x} * \frac{8}{2x+1} = \frac{2(2x+1)}{x} * \frac{8}{2x+1} = \frac{16}{x} \]